ABSOLUTELY SUMMING MULTILINEAR OPERATORS

DAVID PÉREZ-GARCÍA

We give a new definition of absolutely summing multilinear operator and we see that it seems to be the right one, in the sense that we can generalize most of the linear results. In particular, we have multilinear versions of Grothendieck’s Theorem and good relations with some other well known classes of multilinear operators, such as Hilbert-Schmidt, nuclear or integral operators.